# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v2

@@ 4 Experiment
RQ1: Can IAG achieve knowledge augmentation for answering questions over LLMs? (§ 4.4)
RQ2: Does our method have good knowledge activation and generalization abilities? (§ 4.5)
@@
RQ4: What is the role of explicit and implicit imagination modules in IAG and IMcQA? (§ 4.8)
Models Reader Params # Documents NQ TriviaQA WebQ
# closed-book setting
T5-l Roberts et al. (2020a) 770M 0 28.5* 28.7* 30.6*
T5-xl 3b 0 28.30 33.92 34.43
IMcQA-l (Ours) 770M 0 29.32 30.11 32.68
IMcQA-xl (Ours) 3b 0 29.59 35.71 37.40
# Retrieval Augmented Generation
DPR Karpukhin et al. (2020) 110M 100 41.5* 56.8* 41.1*
RAG Lewis et al. (2020) 400M 10 44.5* 56.1* 45.2*
FiD-l Izacard and Grave (2021) 770M 10 46.7* 61.9* 48.1*
FiD-xl 3b 10 50.1* 66.3* 50.8*
# Generation Augmented Generation
GENREAD-l Yu et al. (2023) 770M 10$\dagger$ 40.3* 67.8* 51.5*
GENREAD-xl 3b 10$\dagger$ 42.6* 69.6* 52.6*
# Imagination Augmented Generation (Ours)
IMcQA-l 770M 1$\dagger$ 42.32 65.48 45.28
IMcQA-xl 3b 1$\dagger$ 46.51 68.38 50.45
IMcQA-l 770M 10 49.92 69.67 51.52
IMcQA-xl 3b 5$\ddagger$ 50.87 70.34 52.78
Table 1: QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that only utilize questions; The latter three parts utilize explicit documents. The best results are in bold. * means that those results are from existing papers, $\dagger$ denotes that the documents were generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation). More results can be seen in Appendix A.5.1.
@@ 4.1 Datasets
We evaluate the proposed approach on three public question answering datasets: NaturalQuestions (NQ) Kwiatkowski et al. (2019), WebQuestions (WQ) Berant et al. (2013) and TriviaQA (TQA) Joshi et al. (2017). To evaluate the model performance, we use the exact match (EM) score for evaluating predicted answers Rajpurkar et al. (2016). We provide dataset details in the Appendix A.4.
@@ 4.2 Baselines
For the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and Llama2-13B Touvron et al. (2023) as the basic model. We evaluate with four diverse settings: without retrieval, with retrieval, with LoRA, and using the proposed IMcQA.
@@ 4.3 Implementations
In the pretraining stage, the Imagine Model initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the HyperNetwork, the feedforward neural network (FFN), and the normalization layers Chen et al. (2023). More implementation details and experimental findings are in the Appendix A.3.
@@ 4.4 Main Results
Table 1 shows the performance results, full results including T5-Base are in the Appednidx A.5.1.
As shown in Table 1, when juxtaposed with closed-book models, RAG, and GAG methods, our proposed IAG framework IMcQA method exhibits state-of-the-art performance with the equivalent magnitude of document count and model size.
@@
The following three parts of Table 1 show the experimental results under the open domain setting. Although our method only deals with one short dummy document, it can still achieve results similar to or better than the RAG and GAG methods, which handle 10 documents. The findings reveal that IMcQA exploits imagined condensed text to strike a balance between efficiency and overhead. Moreover, when IMcQA utilizes 10 retrieved documents, it supersedes the performance of FiD and GENREAD with the EM average of +2.26% and +3.06%.
@@ 4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the IMcQA method and the importance of HyperNetwork, we also evaluate its performance in out-of-distribution (OOD) generalizations. Table 2 shows the IID and OOD performance of FiD, and IMcQA methods with different document settings when training on NQ (From NQ generalization to the other two datasets). Full OOD results of three datasets are shown in the Appendix A.5.2.
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, IMcQA can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved ones.
@@
4.6 Zero-shot Results on LLMs
Figure 4 and Figure 5 illustrate the zero-shot results for LLMs implementing IMcQA. This research seeks to explore the possibility of enhancing LLMs via IAG. Due to the high computational demands of training, we only fine-tuned the HyperNetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting. Detailed prompt information can be found in the Appendix A.2.
Models # Docu- NQ
ments NQ TQA WQ
T5 0 22.16 3.18 4.12
IMcQA 0 23.89 6.21 10.94
IMcQA 1$\dagger$ 40.14 46.61 18.92
FiD 10 46.81 53.93 24.02
IMcQA 10 47.01 55.74 24.13
T5-l 0 28.5* 3.18 4.12
IMcQA-l 0 29.32 10.17 14.06
IMcQA-l 1$\dagger$ 42.32 54.80 22.05
FiD-l 10 46.7* 57.93 25.12
IMcQA-l 10 49.92 60.03 25.79
Table 2: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance (IID).
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. IMcQA can enhance knowledge via two main imagination processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. With IAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results. This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs. More results can be seen in A.5.3.
@@ 4.7 Training Cost and Inference Speed-up
Models Training Params # Documents # Avg Tokens Inference Time GPU Memory
T5 220M 0 19.8 79.8s 2828M
IMcQA 139.3M 0 19.8 82.3s 2710M
IMcQA 139.3M 1 522.1 214.6s 2882M
FiD 220M 10 1748.3 683.3s 4358M
GENREAD 220M 10 1912.5 704.8s 4412M
FiD 220M 100 16625.7 1293.2s 19048M
We proceeded to measure the inference speed, documented in GPU time, and training time for 5000 steps on the NQ dataset, using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference.
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared HyperNetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage owing to distillation constraints, our approach achieves efficient reasoning through an extremely lightweight design. Compared to the other two methods, the token processing count is significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets.
@@ 4.8 Ablation Experiment
In this study, we introduced two key imagination processes to stimulate LLMs’ internal knowledge: Explicit Imagination (EI) and Implicit Imagination (II). We particularly examined the influence of different imagination types on performance.
@@
Figure 5: Zero-Shot results (EM, %) of Llama2-13B on three open-domain QA datasets.
The outcomes of Long Context Distillation (LCD) and the application of EI in the HyperNetwork also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains.
@@ 5 Conclusion and Future Work
Appendix A Appendix
A.1 Background
@@
The goal of QA is to learn a distribution function, represented as $p(a|q)$, wherein the models decode a string $a$ that serves as an abstractive answer to a given query $q$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{a}=\arg\max_{a\in\mathcal{V}^{*}}p(a|q,\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to the omission of a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(a|q,c)$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(a|q,c)=\sum\limits^{c_{i}\in c}\limits_{c\in TopK(q)}p(a|q,c_{i})p(c_{i}|q)$, where $Topk(q)$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $q$.
A.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit imagination of the Imagine Model to imagine a short dummy useful document is:
@@ A.3.1 Hyperparameters
LoRA-Base 0$\dagger$ 40000 5e-4 8
IMcQA 0$\dagger$ 50000 1e-3 8
LoRA-l 0$\dagger$ 40000 1e-4 4
IMcQA-l 0$\dagger$ 50000 5e-4 4
LoRA-3b 0$\dagger$ 40000 1e-4 2
IMcQA-3b 0$\dagger$ 50000 1e-4 1
@@
LoRA-Base 1$\dagger$ 37.17 45.20 15.62 19.57 55.37 12.50 14.15 30.89 28.88
IMcQA 1$\dagger$ 40.14 46.61 18.92 24.78 60.75 12.82 17.70 35.24 41.06
@@
LoRA-l 1$\dagger$ 37.61 48.50 20.71 20.54 62.71 14.81 15.36 33.83 39.37
IMcQA-l 1$\dagger$ 42.32 54.80 22.05 26.11 65.48 18.11 18.58 47.46 45.28
@@ A.3.3 Evaluation
A.4 Downstream Evaluation Datasets
@@
  • NaturalQuestions (Kwiatkowski et al. (2019)) contains questions corresponding to Google search queries. The open-domain version of this dataset is obtained by discarding answers with more than 5 tokens, each accompanied by a Wikipedia article containing the answer.
  • TriviaQA (Joshi et al. (2017)) contains questions gathered from trivia and quiz-league websites. The unfiltered version of TriviaQA is used for open-domain question answering, each question is accompanied by pages from web and Wikipedia searches that may contain the answer.
  • WebQuestions (Berant et al. (2013)) contains questions from web queries matched to corresponding entries in FreeBase Bollacker et al. (2008).
A.5 Full Experimental Results
A.5.1 Supervised Performance

### v3

@@ 4 Experiment
RQ1: Can IAG and IMcQA achieve knowledge augmentation for QA over LLMs? (§ 4.4)
RQ2: Does IMcQA have a good out-of-distribution generalization ability? (§ 4.5)
@@
RQ4: What is the role of explicit and implicit imagination modules in IMcQA? (§ 4.7)
@@ 4.1 Datasets
We evaluate the proposed approach on three public question answering datasets: NaturalQuestions (NQ) Kwiatkowski et al. (2019), WebQuestions (WQ) Berant et al. (2013) and TriviaQA (TQA) Joshi et al. (2017). To evaluate the model performance, we use the exact match (EM) score for evaluating predicted answers Rajpurkar et al. (2016). We provide dataset details in the B.4.
@@ 4.2 Baselines
To demonstrate the plug-and-play capability of IMcQA on the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and -13B Touvron et al. (2023) as the basic model. We evaluate with 6 diverse settings: without retrieval, with retrieval, with LoRA, RECITE Sun et al. (2023), HICL Wang et al. (2024) and using the proposed IMcQA.
@@ 4.3 Implementations
In the pretraining stage, the context generator initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the hypernetwork [^5] B.3 contains more implementation and baseline details..
@@ 4.4 Main Results
4.4.1 Supervised Setting
Table 1 shows the performance results, full results including T5-Base are in the C.1. When juxtaposed with closed-book models, RAG, and GAG methods, our proposed IAG framework IMcQA method exhibits state-of-the-art (SOTA) performance with the equivalent magnitude of document numbers.
@@ 4.4.1 Supervised Setting
The following three parts show the experimental results under the open domain setting [^6] Due to memory constraints from online distillation, we only compare the IMcQA-l under the RAG setting using 30 documents with FiD using 100 documents.. Although our method only deals with one short dummy document, it can still achieve results similar to or better than the RAG and GAG methods, which handle 10 documents. The findings reveal that IMcQA exploits imagined condensed text to strike a balance between efficiency and overhead. Beyond this, when IMcQA implements 10 retrieved documents under RAG, it overtakes RFiD’s performance, posting 1.6% higher in NQ, 4.4% in TQA, and 2.7% in WQ. IMcQA employing 10 generated documents under GAG outstrips GENREAD (clustering)’s performance, documenting a 4.5% increment in NQ, 0.7% in TQA, and 1.1% in WQ.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets. The number in parentheses indicates the number of documents used.
4.4.2 Zero-shot Setting
Figure 4 illustrates the zero-shot results for LLMs implementing IMcQA with a frozen Llama2-7B and -13B. This research seeks to explore the possibility of enhancing LLMs via IAG. Due to the high computational demands of training, we only fine-tuned the hypernetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting [^7] Detailed prompts information can be found in the B.2.. More zero-shot setting results can be seen in C.3.
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. IMcQA can enhance knowledge via two main imagination processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. Compared to two other advanced RAG methods, IMcQA using a single document, performs only 1 EM lower than the HICL method Wang et al. (2024) on the TQA but achieves +10% EM on the NQ and +5% EM on the WQ. With IAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results (Figure 6). This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs.
Models # Docs Base (220M) Large (800M)
NQ TQA WQ NQ TQA WQ
T5 0 22.16 3.18 4.12 28.5* 3.18 4.12
IMcQA 0 23.89 6.21 10.94 29.32 10.17 14.06
FiD 10 46.81 53.93 24.02 46.7* 57.93 25.12
IMcQA 10 47.01 55.74 24.13 49.92 60.03 25.79
LoRA 1 5.43 3.15 4.02 4.42 6.50 7.87
IMcQA 1 40.14 46.61 18.92 42.32 54.80 22.05
Table 2: IID and OOD results. The performance on three open-domain datasets for the model trained on NQ is reported, with the underlined values indicating the IID performance. Full OOD results and details of three datasets are shown in the C.2.
@@ 4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the IMcQA method and the importance of hypernetwork, we also evaluate its performance in OOD generalizations. Table 2 shows the IID and OOD performance of FiD, and IMcQA methods with different document settings when training on NQ (From NQ generalization to the other two datasets).
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, IMcQA can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved documents.
@@
Models Training Params # Documents # Avg Tokens Inference Time GPU Memory Training Time
T5 220M 0 19.8 79.8s 2828M 0.9h
IMcQA 139.3M 0 19.8 82.3s 2710M 1.2h
IMcQA 139.3M 1 522.1 214.6s 2882M 1.7h
FiD 220M 10 1748.3 683.3s 4358M 2.3h
GENREAD 220M 10 1912.5 704.8s 4412M -
FiD 220M 100 16625.7 1293.2s 19048M 5.8h
@@ 4.6 Training Cost and Inference Speed-up
We proceeded to measure the inference speed documented in GPU time and training time for 5000 steps on the NQ dataset using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference.
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared hypernetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage due to distillation constraints, IMcQA achieves efficient reasoning through an extremely lightweight design, saving more than half the training time compared to methods using a large number of documents (×$\times 0.3$). Compared to the other two methods, the processed tokens are significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
@@ 4.7 Ablation Experiment
This study introduces two key imagination processes to stimulate LLMs’ internal knowledge: Explicit Imagination (EI) and Implicit Imagination (II). We particularly examined the influence of different imagination types on performance.
@@
The outcomes of Long Context Distillation (LCD) and the application of EI in the hypernetwork also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains. The experiments and analysis above demonstrate the importance of each component and the effectiveness of our IMcQA method.
@@ 5 Conclusion and Future Work
Appendix A Method
A.1 Comparison of Three Paradigms
Compared to RAG and GAG, our method has certain limitations, such as requiring a more complex training process and the necessity of training a model. Similar to the GAG method, which uses a master’s degree in law as a knowledge base, our method also struggles to generate content when encountering new and unknown world knowledge, which presents a challenge that needs to be addressed. Additionally, the knowledge base might be affected by knowledge gaps in low-resource settings where there is a lack of a comprehensive knowledge base.
Next, we compare IAG, RAG, and GAG across four criteria for a more intuitive understanding. From the table 4, it can be observed that the document relevance obtained by IAG and GAG is higher, while RAG heavily relies on the retriever and external knowledge base. In terms of document length usage, IAG only needs to use a virtual document, greatly reducing the number of tokens. Therefore, IAG is superior to the other two methods in terms of reasoning time.
A.2 Context Generator
There are two main goals in the pre-training of the model in the first stage of IAG (context generator): first, to improve its ability as a document generator by learning to generate rich and concise documents; second, to introduce some external knowledge that the model initially does not possess. It’s worth noting that the second goal is crucial, as the model may encounter knowledge it has not yet learned. Thus, IAG does not rely on external large models or retrievers for external reasoning and can complete reasoning independently.
A.3 Hypernetwork
Hypernetworks have gained significant attention in recent years due to their potential to enhance various aspects of neural network performance. In this section, we analyze the reasons for employing hypernetworks in detail:
Hypernetworks, which are neural networks designed to generate the weights of other networks, allow for dynamic adjustment of model parameters. This adaptability enables the model to better suit different tasks and datasets, thereby improving overall performance. By utilizing hypernetworks, the number of models that need to be trained individually can be significantly reduced. Traditional methods require separate models for each task, whereas hypernetworks can generate weights for multiple tasks. This capability enhances training efficiency. In our task, we use hypernetworks to generate adapters for the question and input, which are then inserted into the model. This helps the model incorporate the knowledge targeted by the question, corresponding to implicit imagination. Compared to traditional efficient fine-tuning, this process is more aligned with the goal of imagination.
Hypernetworks can capture the commonalities and differences between various tasks by learning to generate weights. This ability to generalize across tasks improves the model’s performance on unseen data, making it more robust in diverse scenarios. In multi-task learning or meta-learning scenarios, hypernetworks can considerably reduce the need for storing multiple independent models. A hypernetwork only needs to store a single generating network and some shared parameters, thus significantly decreasing the storage space required. Hypernetworks can quickly generate new weights to adapt to new tasks as they arise. This rapid adaptation capability is particularly useful in applications that require frequent updates or expansions. In our experiments 4.5, we also found that using a hypernetwork can significantly enhance the generalization ability for tasks. This is because it not only retains knowledge within the domain-specific modules but also learns to generate question-targeted knowledge to be inserted into the model.
Appendix B Experimantal Settings
B.1 Background
@@
The goal of QA is to learn a distribution function, represented as $p(a|q)$, wherein the models decode a string $a$ that serves as an abstractive answer to a given query $q$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{a}=\arg\max_{a\in\mathcal{V}^{*}}p(a|q,\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to omitting a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(a|q,c)$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(a|q,c)=\sum\limits^{c_{i}\in c}\limits_{c\in TopK(q)}p(a|q,c_{i})p(c_{i}|q)$, where $Topk(q)$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $q$.
B.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit imagination of the context generator to imagine a short dummy useful document is:
@@ B.2 Prompts for Explicit Imagine with LLMs
LoRA-Base 0 40000 5e-4 8
IMcQA 0 50000 1e-3 8
LoRA-l 0 40000 1e-4 4
IMcQA-l 0 50000 5e-4 4
LoRA-3b 0 40000 1e-4 2
IMcQA-3b 0 50000 1e-4 1
@@
LoRA-Base 1 37.17 45.20 15.62 19.57 55.37 12.50 14.15 30.89 28.88
IMcQA 1 40.14 46.61 18.92 24.78 60.75 12.82 17.70 35.24 41.06
@@
LoRA-l 1 37.61 48.50 20.71 20.54 62.71 14.81 15.36 33.83 39.37
IMcQA-l 1 42.32 54.80 22.05 26.11 65.48 18.11 18.58 47.46 45.28
@@ B.3.2 Baselines
RFiD   Wang et al. (2023a) uses the encoder of FiD to distinguish between causal and incidental features, and guides the decoder to generate answers based on this distinction.
EAR   Chuang et al. (2023) significantly enhances the traditional sparse retrieval method BM25 by connecting query expansion models and retrievers.
FILCO   Wang et al. (2023d) identifies useful context based on lexical and information-theoretic methods.
@@ B.3.3 Evaluation
Dataset Train Dev Test
WebQ 3,417 361 2,032
NQ 79,168 8,757 3,610
TQA 78,785 8,837 11,313
Table 7: Open-Domain QA dataset statistics
B.4 Downstream Evaluation Datasets
@@
  • NaturalQuestions Kwiatkowski et al. (2019) contains questions corresponding to Google search queries. The open-domain version of this dataset is obtained by discarding answers with more than 5 tokens, each accompanied by a Wikipedia article containing the answer.
  • TriviaQA Joshi et al. (2017) contains questions gathered from trivia and quiz-league websites. The unfiltered version of TriviaQA is used for open-domain question answering, each question is accompanied by pages from web and Wikipedia searches that may contain the answer.
  • WebQuestions Berant et al. (2013) contains questions from web queries matched to corresponding entries in FreeBase Bollacker et al. (2008).
Table 7 presents detailed statistics of the dataset sizes, including the training, development, and test sets. We note that all our models are trained exclusively on the training data, and we did not include the development data in our training process. Therefore, the performance numbers reported in the paper for the dev and test data are independent of the training data.
Figure 6: Zero-Shot results (EM, %) of Llama2-13B on three open-domain QA datasets. The number in parentheses indicates the number of documents used.
Appendix C Full Experimental Results
C.1 Supervised Performance